The supply curve for sugar is given as Supply: Qs = 1750 + 24P and demand for sugar is given as Demand: Ds = 3270 – 26P; where P is price in SR per kg. What will be the equilibrium price and equilibrium quantity in the free market? 


1
Expert's answer
2021-03-11T07:23:31-0500

Market equilibrium occurs when "Q_d=Q_s":


"3270-26P_E=1750+24P_E,""1520=50P_E,""P_E=SR\ 30.4."

Finally, we can find the equilibrium quantity:


"Q_E=3270- 26\cdot SR\ 30.4 =2479.6."